OpenStudy (anonymous):

Solve the system of equations using the substitution method. Enter the answer as an ordered pair. 3x - y = 3 y = 2x + 2

OpenStudy (vijay):

-1, 0

OpenStudy (anonymous):

3x - (2x +2) = 3, now you continue...

OpenStudy (anonymous):

y= -3x +y

OpenStudy (anonymous):

3x-(2x-2)=3 3x-2x-2=3 x-2=3 x=5 i think...my brain is kinda fryed from all the math lol

OpenStudy (anonymous):

substitution= get the equation in the term of one variable, and then plug that into the second equation. ex. a+b=1, and a*b=0 Get the 1st equation in terms of a (or b) a=1-b plug that in wherever u see a in the second equation. (1-b)*b=1

OpenStudy (anonymous):

For this one, I'd get the 1st equation in terms of y 3x-y=3 -y=3-3x y=-3+3x

OpenStudy (anonymous):

x = 5 y = 12

OpenStudy (anonymous):

as an ordered pair that would (5,12)

OpenStudy (anonymous):

Now plug that in wherever u see y in the second equation. -3+3x=2x+2 x=5

OpenStudy (anonymous):

Now that u know x=5, go back to one of the original equations, plug in a 5 wherever u see x, and solve for y

OpenStudy (anonymous):

y=2(5)+2 y=10+2 y=12

OpenStudy (vijay):

many answers r suitable for this questions

OpenStudy (anonymous):

So x=5, y=12.